# Decompositional Reasoning for Graph Retrieval with Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) excel at many NLP tasks, but struggle with multi-hop reasoning and factual consistency, limiting their effectiveness on knowledge-intensive tasks like complex question answering (QA). Linking Knowledge Graphs (KG) and LLMs has shown promising results, but LLMs generally lack the ability to reason efficiently over graph-structured information. To tackle this problem, we propose a novel retrieval approach that integrates textual knowledge graphs into the LLM reasoning process via query decomposition. Our method decomposes complex questions into sub-questions, retrieves relevant textual subgraphs, and composes a question-specific knowledge graph to guide answer generation. For that, we use a weighted similarity function that focuses on both the complex question and the generated subquestions to extract a relevant subgraph, which allows efficient and precise retrieval for complex questions and improves the performance of LLMs on multi-hop QA tasks. This structured reasoning pipeline enhances factual grounding and interpretability while leveraging the generative strengths of LLMs. We evaluate our method on standard multi-hop QA benchmarks and show that it achieves comparable or superior performance to competitive existing methods, using smaller models and fewer LLM calls. Source code will be available upon acceptance.
Decompositional Reasoning for Graph Retrieval with Large Language Models
@@ 1 Introduction
Figure 1: Illustration of our decompositional retrieval-based reasoning method. Our method decomposes the question into sub-questions, performs iterative, context-aware retrieval conditioned on previous answers, and merges the resulting subgraphs for guided reasoning.
@@
In this work, we address these issues by proposing a method for knowledge-guided decompositional reasoning with LLMs. Our method injects structured knowledge into the reasoning process using textualized knowledge graphs. Specifically, we decompose complex questions into sub-questions, retrieve relevant subgraphs from a textual knowledge graph, and merge them for structured and reasoning-enhanced retrieval (Figure 1). The obtained graph is then used to guide LLMs toward generating more accurate and interpretable answers. This hybrid approach enhances both the factual correctness and the transparency of LLM predictions, particularly in settings requiring multi-step reasoning over domain knowledge. To verify the effectiveness of our approach, we conduct extensive experiments on benchmark datasets for complex question answering, namely CWQ (Talmor and Berant, 2018) and WebQSP (Yih et al., 2016). We compare our method against standard prompting techniques, as well as existing state-of-the-art approaches that combine LLMs with knowledge graphs. Results show that our method achieves consistent improvements in accuracy without increasing the number of parameters for the LLM or the number of LLM calls, which shows the efficiency of our method. Our contributions are as follows:
  • We develop a novel knowledge graph retrieval method that uses query decomposition, helping the LLM to reason over structured data for complex questions.
  • We introduce a hybrid similarity function that uses the complex question and its decomposition to guide the retrieval process.
  • We demonstrate improvements in accuracy and factual consistency on multi-hop QA benchmarks.
  • Our method reduces the number of LLM-calls compared to other baselines, achieving a 3$\times$ to 5$\times$ reduction for both datasets.
@@ 2.2 LLMs and graphs
Figure 2: Overview of our retrieval method: the complex question is first decomposed into smaller subquestions, for which we iteratively perform retrieval and answer generation; once the retrieval is done for all subquestions, we merge the subgraphs and give the result as a hard (textualized graph) and a soft prompt (graph encoder output) to the model.
@@
LLMs and graph neural networks (GNNs) can also be used together (Xu et al., 2024) (He et al., 2024a), each complementing the other. Graphs can be used to inject knowledge into LLMs via methods like structured prompting (Baek et al., 2023) (Zhang et al., 2024) or retrieval-based augmentation (Lewis et al., 2020) (Peng et al., 2024). LLMs can support and enhance graph-centred tasks (Pan et al., 2024) by performing entity linking, relation extraction, or even link prediction (Shu et al., 2025), which largely improves the graph’s coverage. LLMs have also been explored as generators of graph-structured outputs or as interpretable reasoning agents over graphs using intermediate symbolic steps. In such hybrid frameworks, LLMs benefit from the structure and factual reliability of graphs, while graphs gain from the generalization and language understanding ability of LLMs (Pan et al., 2024). Nonetheless, most existing methods remain heuristic and lack a principled understanding of how best to align symbolic and neural representations (Cheng et al., 2025).
@@ 3 Related Work
Different methods have already demonstrated promising results on Knowledge Graph Question Answering (KGQA) tasks. He et al. (2024b) retrieves a subgraph from a textual knowledge graph and feeds it to the LLM without any explicit reasoning step, which can hinder performance on complex questions. Other existing techniques introduce some reasoning mechanisms within their framework: Sun et al. (2024) performs iterative entity and relation explorations, and directly reasons on the obtained paths. Similarly, Chen et al. (2024) uses task decomposition and then performs multiple cycles of exploration, memory updating, and evaluation. Performing iterative calls to the LLM has many advantages, but both mentioned techniques require using a relatively large model (LLaMa-2-70B, GPT-3.5 / GPT-4…) for planning and evaluation. In contrast, our method focuses on retrieving a more pertinent graph rather than answering iteratively, and uses fewer LLM calls—which can be controlled by the number of generated subquestions. Other methods like Luo et al. (2024) first predict reasoning paths as plans and search for those paths within the knowledge graph. Given that the LLM does not have any prior knowledge of the relations within the knowledge graph, the technique requires knowledge distillation into the LLM to generate faithful relation paths. Our method does not require any fine-tuning on the LLM, which reduces the cost of usage and the preprocessing time for new datasets.
@@ 4 Method
The overall pipeline for our method is presented in Figure 2. In order to tackle complex questions, we first decompose a complex question into a set of logically ordered subquestions. We then perform an iterative retrieval cycle by performing retrieval on the graph for each subquestion that we obtain. The results for the multiple retrievals are then combined into a single graph, which is then used to generate the final answer to the complex question.
@@ 4.2 Hybrid Entity Retrieval
For the retrieval part of the generation pipeline, we want to obtain a subgraph for each subquestion. Considering each subquestion independently might lead to very distinct subgraphs; moreover, the subquestions can lack sufficient contextual information on their own to retrieve all the relevant nodes and edges from the knowledge graph. To address this, we introduce a hybrid retrieval method that combines both the subquestion and the original complex question, allowing the model to benefit from the specificity of the former and retain the broader context provided by the latter. Our hybrid retrieval mechanism is implemented through a weighted similarity function, controlled by a parameter $\alpha$, which balances the influence of both components. Figure 3 presents the equations for both node and edge retrieval.
When performing retrieval on the graph for the subquestion $q_{i}$, we keep track of the previous answer (answer $a_{i-1}$ to subquestion $q_{i-1}$). This is crucial, as the answer to $q_{i}$ might depend on the answer to $q_{i-1}$. Before retrieval, we embedded our complex question, the subquestions, and the textual attributes of the nodes/edges in the graph using a Sentence Transformer embedding model (see Appendix B for details). After having retrieved all necessary nodes and edges, we build a connected subgraph from these elements, following work done in He et al. (2024b). The connectivity of the graph is enforced by the Prize-Collecting Steiner Tree (PCST) algorithm (Bienstock et al., 1993), which optimizes the selection of a subgraph of maximum value based on node/edge weights and query similarity, under a size constraint.
${V_{k}}^{i}=\operatorname*{argtopk}_{n\in V}\left[\alpha\,\cos(z_{i},z_{n})+(1-\alpha)\,\cos(z_{Q},z_{n})\right]$
${E_{k}}^{i}=\operatorname*{argtopk}_{e\in E}\left[\alpha\,\cos(z_{i},z_{e})+(1-\alpha)\,\cos(z_{Q},z_{e})\right]$
Figure 3: Hybrid retrieval: ${V_{k}}^{i}$ and ${E_{k}}^{i}$ denote the top nodes and edges from $G=(V,E)$, ranked by a weighted similarity to the subquestion $q_{i}$ and original question $Q$. $z_{i}$, $z_{Q}$, $z_{n}$, and $z_{e}$ are the embeddings of $q_{i}$, $Q$, node $n$, and edge $e$, respectively.
@@ 4.3 Subgraphs Merging
After retrieving subgraphs corresponding to each subquestion, we proceed to merge them in order to link relevant information and remove redundancy. Each subgraph is initially connected, as it is constructed using the PCST algorithm. To form the final graph, we take the union of all distinct nodes and edges across all subgraphs. Importantly, we do not directly enforce full connectivity during this merging step, as doing so would require introducing virtual edges, which could potentially compromise the semantic integrity of the graph or resort to computationally expensive graph expansion methods.
@@ 4.4 Answer Generation
Once we obtained the merged graph for the complex question, we pass it to our LLM in two different ways, following the generation process described in He et al. (2024b): we provide a textualized version of the graph in the prompt, and also pass the graph through a trained graph encoder (Shi et al., 2021) followed by a linear projection layer. Providing the encoded graph as a soft prompt guides the LLM’s response by feeding a trained embedding vector to the self-attention layers of the language model. When answering the complex question, we include the merged graph and its textual description in the prompt; we chose not to include the answers to the subquestions in the final prompt, as a single prior error can force the model to give a wrong answer, even when the graph contains the correct answer.
@@ 5.1 Benchmarks
We evaluate our method on two different Question-Answering (QA) benchmarks to assess the quality of our results: CWQ (ComplexWebQuestions) (Yih et al., 2016) and WebQSP (WebQuestionsSemanticParses) (Talmor and Berant, 2018), which are both based on the Freebase (Bollacker et al., 2008) knowledge base. CWQ is a complex QA benchmark that focuses on multi-hop questions. As it needs the integration of multiple facts, it benefits from compositional reasoning, making it a suitable benchmark for our approach. WebQSP, on the other hand, contains a wide range of simple and factual questions. It also includes SPARQL annotations that we do not use in this work. We use the preprocessed version of the dataset provided in Luo et al. (2024).
@@ 5.3 Choice of language models
Our method requires two distinct capabilities, each handled by different classes of models. First, strong decompositional reasoning is needed to break down the complex question into logically ordered, comprehensive, and atomic subquestions. There, we use a Qwen-32B model distilled from Deepseek-R1 (DeepSeek-AI et al., 2025) for its advanced reasoning abilities. Second, we need efficient models to answer the subquestions and generate the final answer. For this, we experiment both with LLaMA-2-7B and LLaMA-2-13B (Touvron et al., 2023b). We also propose a "Hybrid 7B/13B" setting in which the 7B model answers the subquestions, while the 13B model handles the final complex question answer. The rationale is that atomic subquestions are simple and can be handled by a smaller model, while the final answer—requiring the integration of the full merged graph—benefits from the greater capacity of a larger model. This setting leverages model efficiency by allocating larger capacity only where necessary. We evaluate both uniform and hybrid settings in Section 6.
5.4 Balancing the Retrieval Query
Using only an atomic subquestion for retrieval can lead to ineffective results, as it lacks the broader context of the original complex question. To address this, we propose balancing the influence of the complex question and the current subquestion in the retrieval query embedding. We introduce an $\alpha$ parameter (Section 4.2) that controls this trade-off via a weighted average of their respective query embeddings. As shown in Figure 3, $\alpha$ determines the contribution of each: lower values (close to 0) emphasize the subquestion, while higher values shift focus toward the original complex question. When $\alpha=1$, retrieval is based solely on the complex question, without any decompositional reasoning, as in He et al. (2024b) (see Figure 4).
@@ 6.1 Influence of $\alpha$ parameter
During retrieval, we use both the complex question and its subquestions, with the $\alpha$ parameter controlling their relative importance in the query (Figure 3). We vary $\alpha$ and report model accuracy in Figure 4. We observe that using a larger model (13B) in the final answer stage (7B/13B and 13B setups) significantly outperforms the 7B-only setup, however using such a model for answering the subquestions offers no clear benefit, as we see, the hybrid 7B/13B and 13B-only setups yield similar results. Across all setups, extreme $\alpha$ values (near 0 or 1) underperform, and intermediate values work best. This supports the need to balance focus between subquestions and the main question during retrieval. In the rest of the paper, we use $\alpha=0.7$.
Varying $\alpha$ also impacts the structure of the retrieved graph, potentially affecting the connectivity constraint previously ensured for subgraphs by the PCST algorithm. Higher $\alpha$ leads to more connected and denser merged graphs (Figures 9, 10), while lower values produce more distinct subgraphs and a sparser, occasionally disconnected graph. Although we observe connected graphs empirically yield better performance (Figure 5), disconnected ones remain rare in proportion (Figure 9). Despite the drop in performance in these cases, results remain competitive with state-of-the-art methods (Table 1). We discuss the statistical significance of these results in Appendix A.
@@
Since our focus is on improving retrieval, we report the Exact Matching scores for different $\alpha$ values. Exact Matching score is defined as the percentage of graphs containing a node that exactly matches the answer label. We observe in Figure 6 that focusing on the subquestions rather leads to Exact Matching scores. Setting $\alpha=1$ serves as a sanity check to verify that we obtain similar metrics to He et al. (2024b). Additional results for Exact Matching can be found in Appendix A. We observe similar results for the Matching score: we define this metric as the percentage of retrieved graphs that contain a node very similar to the answer label (based on a cosine similarity between embeddings, using a similarity threshold of 0.9). This more flexible metric allows to check the presence of highly related nodes in the retrieved graph.

### v3

# DAGR: Decomposition Augmented Graph Retrieval with LLMs

@@ Abstract
Large Language Models (LLMs) excel at many Natural Language Processing (NLP) tasks, but struggle with multi-hop reasoning and factual consistency, limiting their effectiveness on knowledge-intensive tasks like complex question answering (QA). Linking Knowledge Graphs (KG) and LLMs has shown promising results, but LLMs generally lack the ability to reason efficiently over graph-structured information. To address this challenge, we introduce DAGR, a retrieval method that leverages both complex questions and their decomposition in subquestions to extract relevant, linked textual subgraphs. DAGR first breaks down complex queries, retrieves subgraphs guided by a weighted similarity function over both the original and decomposed queries, and creates a question-specific knowledge graph to guide answer generation. The resulting Graph-RAG pipeline is suited to handle complex multi-hop questions and effectively reason over graph-structured data. We evaluate DAGR on standard multi-hop QA benchmarks and show that it achieves comparable or superior performance to competitive existing methods, using smaller models and fewer LLM calls. Source code will be available upon acceptance.
DAGR: Decomposition Augmented Graph Retrieval with LLMs
@@ 1 Introduction
Figure 1: Illustration of DAGR, our novel decompositional retrieval method. We first decompose the complex question into sub-questions, perform iterative, context-aware retrieval conditioned on previous answers, and merges the resulting subgraphs. The resulting graph can then be used to ground the generation of an LLM.
@@
In this work, we propose DAGR, a retrieval method designed to improve structured knowledge access for LLMs through decompositional reasoning over textualized knowledge graphs. Given a complex question and a knowledge graph, DAGR decomposes it into sub-questions, retrieves relevant subgraphs using a hybrid similarity function informed by both the original and decomposed queries, and merges them into a coherent, question-specific textual graph. As seen on Figure 1, the resulting structured graphs can be plugged into Graph-RAG pipelines, enhancing their ability to guide LLMs toward more factual and interpretable answers. This novel retrieval method allows precise and coherent multi-step retrieval for complex questions, and removes the need of using very large models or fine-tuning the LLM on a specific graph. The hybrid similarity function fuses the richness of the generated subquestions as well as the anchoring of the initial complex question, helping to retrieve multiple coherent and linked subgraphs. We validate DAGR on the CWQ (Talmor and Berant, 2018) and WebQSP (Yih et al., 2016) multi-hop QA datasets, demonstrating that our decompositional retrieval method achieves state-of-the-art performance when paired with smaller, frozen LLMs, while requiring significantly fewer LLM calls.
  • We propose DAGR, a retrieval method that augments the complex query with its decomposition to retrieve relevant linked subgraphs.
  • We introduce a hybrid similarity function that integrates the full query and its decomposition to guide subgraph selection.
  • We obtain state-of-the-art results on multi-hop QA benchmarks, without using large or fine-tuned LLMs.
[…]
@@ 2.2 LLMs and graphs
Figure 2: Usage of our decompositional retrieval method: 1) The complex question is first decomposed into atomic subquestions ; 2) We conditionally perform retrieval and answer generation; once the retrieval is done for all subquestions, we merge the subgraphs ; 3) The resulting graph is given as a hard (textualized graph) and soft prompt (graph encoder output) to the model.
@@
LLMs and graph neural networks (GNNs) can also be used together (Xu et al., 2024) (He et al., 2024a), each complementing the other. Graphs can be used to inject knowledge into LLMs via methods like structured prompting (Baek et al., 2023) (Zhang et al., 2024a) or retrieval-based augmentation (Lewis et al., 2020) (Peng et al., 2024). LLMs can support and enhance graph-centred tasks (Pan et al., 2024) by performing entity linking, relation extraction, or even link prediction (Shu et al., 2025), which largely improves the graph’s coverage. LLMs have also been explored as generators of graph-structured outputs or as interpretable reasoning agents over graphs using intermediate symbolic steps. In such hybrid frameworks, LLMs benefit from the structure and factual reliability of graphs, while graphs gain from the generalization and language understanding ability of LLMs (Pan et al., 2024). Nonetheless, most existing methods remain heuristic and lack a principled understanding of how best to align symbolic and neural representations (Cheng et al., 2025).
@@ 3 Related Work
Several methods have demonstrated strong performance on Knowledge Graph Question Answering (KGQA) tasks by combining LLMs with structured information. He et al. (2024b) retrieves a single subgraph from a textual knowledge graph and directly feeds it into an LLM for answer generation. However, the lack of an explicit reasoning step limits its effectiveness on complex, multi-hop questions, where structured reasoning is essential. Other approaches incorporate reasoning more explicitly: for instance, Sun et al. (2024) introduces iterative entity and relation exploration guided by the LLM, while Chen et al. (2024) decomposes tasks and performs multiple reasoning cycles involving exploration, memory updates, and evaluation. While these iterative strategies improve reasoning quality, they rely on a high number of LLM calls and require large models (e.g., LLaMA-2-70B, GPT-3.5, or GPT-4) to plan and evaluate effectively, which increases the computational cost and inference time. Luo et al. (2024) instead predicts relation paths as intermediate plans, but requires training or distilling structured knowledge into the LLM so it can generate faithful and executable relation paths. In contrast, our method DAGR focuses on efficiently constructing question-specific subgraphs through decompositional retrieval, instead of relying on large models, dataset-specific fine-tuning, or LLM-guided exploration. DAGR breaks down complex questions into sub-questions and retrieves linked subgraphs using a hybrid similarity function, composing a targeted knowledge graph to support answer generation. This approach enables strong performance on complex QA tasks while using smaller, frozen LLMs, and reduces the number of LLM calls. Additionally, the lack of LLM fine-tuning makes our method DAGR adaptable to new datasets with minimal overhead.
@@ 4 Method
The overall generation pipeline that uses our method is presented in Figure 2. In order to tackle complex questions, we first decompose a complex question into a set of logically ordered subquestions. We then perform an iterative retrieval cycle by performing retrieval on the graph for each generated subquestion. The obtained subgraphs are then merged into a single knowledge graph. For the answer generation, the obtained graph is then fed to the LLM, following work done by He et al. (2024b).
@@ 4.2 Hybrid Entity Retrieval
For each generated subquestion $q$, we want to obtain a subgraph $G_{q}$. Performing a retrieval step for each subquestion helps to exploit the richness of the decomposition, using all implicit reasoning steps to answer the complex question $Q$. But treating each subquestion independently might lead to very distant subgraphs; moreover, the subquestions can lack sufficient contextual information on their own to retrieve all the relevant nodes and edges from the knowledge graph. To address this issue, we introduce a hybrid similarity function (Figure 3) that combines both the subquestion and the original complex question, allowing the model to benefit from the specificity of $q$ and retain the broader context provided by the latter $Q$. In our hybrid similarity function, the influence of both components is controlled by a parameter $\alpha$.
$\underset{\alpha}{\text{DAGR}}(z,z_{q},z_{Q})=\alpha\,\cos(z,z_{q})+(1-\alpha)\,\cos(z,z_{Q})$
${V_{k}}^{q}=\operatorname*{argtopk}_{n\in V}\underset{\alpha}{\text{DAGR}}(z_{n},z_{q},z_{Q})$
${E_{k}}^{q}=\operatorname*{argtopk}_{e\in E}\underset{\alpha}{\text{DAGR}}(z_{e},z_{q},z_{Q})$
Figure 3: Hybrid graph-based retrieval using our proposed similarity function (with parameter $\alpha$), which combines subquestion-level and global question-level semantics. The sets ${V_{k}}^{q}$ and ${E_{k}}^{q}$ represent the top-$k$ nodes and edges in the graph $G=(V,E)$, retrieved based on their similarity to subquestion embedding $z_{q}$ and the original question embedding $z_{Q}$.
Before retrieval, we embedded our different components (complex question, the subquestions, and the textual attributes of the nodes/edges in the graph) using a Sentence Transformer embedding model (see Appendix B for details). When performing retrieval on the graph for the subquestion $q_{i}$, we keep track of the answer $a_{i-1}$ to the previous subquestion $q_{i-1}$. This is crucial, as the answer to $q_{i}$ usually depends on the answer to $q_{i-1}$. Therefore, we combine the subquestion $q_{i}$ and the previous answer $a_{i-1}$ in a single embedding $z_{q}=z(a_{i-1},q_{i})$. For simplicity, we choose to concatenate the previous answer element and the next subquestion, allowing context sharing from the previous step. After having retrieved all necessary nodes and edges, we build a connected subgraph from these elements, ensuring that a subgraph represents a single component with linked graph entities. The connectivity of the graph is enforced by the Prize-Collecting Steiner Tree (PCST) algorithm (Bienstock et al., 1993), which optimizes the selection of a subgraph of maximum value based on node/edge weights and query similarity, under a size constraint. This method also allows us to control the size of the retrieved subgraph, as a large subgraph might not always fit in the context size of the model. Impact of the chosen maximal size of a subgraph is studied in later analysis.
@@ 4.3 Subgraphs Merging
After retrieving subgraphs corresponding to each subquestion, we proceed to merge them in order to link relevant information and remove redundancy. To form the final graph, we take the union of all distinct nodes and edges across all subgraphs, removing duplicate entities: $G^{*}=\bigcup_{i=1}^{n}G_{i}$. A crucial observation is that the resulting merged graph is almost always connected, thanks to the anchoring used in the similarity function at each step, by using the initial complex question $Q$. This allows us to bypass enforcing connectivity by introducing virtual edges, which could potentially compromise the semantic integrity of the graph or resort to computationally expensive graph expansion methods.
@@ 4.4 Answer Generation
Once we obtain the merged graph $G^{*}$ from the different subgraphs, we pass it to the LLM, following the generation process described in He et al. (2024b): we provide a textualized version of the graph in the prompt, and also pass the graph through a trained graph encoder (Shi et al., 2021) followed by a linear projection layer. Providing the encoded graph as a soft prompt guides the LLM’s response by feeding a trained embedding vector to the self-attention layers of the language model. When answering the complex question, we chose not to include the answers to the subquestions in the final prompt, as a single prior error can force the model to give a wrong answer. Instead, the prompt only contains $Q$ and merged graph.
@@ 5.1 Benchmarks
We evaluate our method on two different Question-Answering (QA) benchmarks to assess the quality of our results: CWQ (ComplexWebQuestions) (Yih et al., 2016) and WebQSP (WebQuestionsSemanticParses) (Talmor and Berant, 2018), which are both based on the Freebase (Bollacker et al., 2008a) knowledge base. CWQ is a complex QA benchmark that focuses on multi-hop questions. As it needs the integration of multiple facts, it benefits from compositional reasoning, making it a suitable benchmark for our approach. WebQSP, on the other hand, contains a wide range of simple and factual questions. It also includes SPARQL annotations that we do not use in this work. This benchmark is used to evaluate the performance of our method on simpler questions, that may not always require decomposition. We use the preprocessed version of the dataset provided in Luo et al. (2024).
@@ 5.3 Choice of language models
Our method requires the usage of two language models, each having a different focus. First, strong decompositional reasoning is needed to break down the complex question into logically ordered, comprehensive, and atomic subquestions. We use a Qwen-32B model distilled from Deepseek-R1 (DeepSeek-AI et al., 2025) for its advanced reasoning abilities. We test different model sizes at the decomposition step (7B, 14B and 32B) and we evaluate the quality of the generated subquestions (see Table 4 and Table 5 in Appendix A), and we show that the 32B model produces superior decomposition compared to other models. We do not consider larger models, as the model size would become a strong usage limitation. Second, we need an efficient model to iteratively answer the subquestions and generate the final answer. For this, we experiment both with LLaMA-2-7B and LLaMA-2-13B (Touvron et al., 2023b). We also propose a "Hybrid 7B/13B" setting in which the 7B model answers the subquestions, while the 13B model handles the final complex question. The rationale is that atomic subquestions are simple and can be handled by a smaller model, while the final answer, requiring the integration of the full merged graph, benefits from the greater capacity of a larger model. This setting leverages model efficiency by allocating larger capacity only where necessary. We evaluate both uniform and hybrid settings in Section 6.
5.4 Balancing the Hybrid Retrieval
Using the subquestion alone for retrieval can lead to ineffective results, as it lacks the broader context of the original question. To address this, we balance the influence of the complex question and the current subquestion in the retrieval query. The $\alpha$ parameter (Section 4.2) controls this trade-off via a weighted average of their respective query embeddings. As shown in Figure 3, $\alpha$ determines the contribution of each: lower values emphasize the subquestion, while higher values shift focus toward the original complex question. When $\alpha=1$, retrieval is based solely on the complex question, without any decompositional reasoning, as in He et al. (2024b). We experiment with different values of $\alpha$ to decide on the importance that both elements should have during the retrieval process.
@@ 6.1 Influence of $\alpha$ parameter
During retrieval, we use both the complex question and the corresponding subquestions, with the $\alpha$ parameter controlling their relative importance in the query (Figure 3). We vary $\alpha$ and report model accuracy in Figure 4.
The first observation is that the value of $\alpha$ has a strong impact on the resulting graph topology, affecting the connectivity and density of the resulting merged graph. Experiments show that higher $\alpha$ leads to more connected and denser graphs (Figures 5, 9), while lower values produce more distinct subgraphs and a sparser, occasionally disconnected graph. The topology of the graphs is highly important for the LLM that will use the retrieved graph: we show (Appendix A) that connected graphs empirically yield better performance, although disconnected graphs remain rare in proportion, as seen on Figure 5. We also discuss the statistical significance of these results in Appendix A.
The second observation is that the $\alpha$ parameter influences the precision (Matching and Exact Matching) of the retrieved entities. Exact Matching score is defined as the percentage of graphs containing a node that exactly matches the answer label (or how often a graph contains the answer to the question asked. We define Matching as the percentage of retrieved graphs that contain a semantically close node to the answer label (using standard cosine similarity between embeddings, and a similarity threshold of 0.9). This metric is more flexible and helps checking the presence of highly related nodes in the retrieved graph. Here, we report the Exact Matching score for different $\alpha$ values, and we observe in Figure 6 that focusing on the subquestions leads to a higher Exact Matching score. Naturally, by setting $\alpha=1$, we obtain similar metrics to He et al. (2024b). We observe similar results for the Additional similar results for Matching and Exact Matching (using different graph size) can be found in Appendix A.
@@
In the end, we observe that using a larger model (13B) in the final answer stage (7B/13B and 13B setups) significantly outperforms the 7B-only setup ; however, using the 13-B model for iterative subquestions answering offers no clear benefit. We observe indeed that the hybrid 7B/13B and 13B-only setups yield similar results. Across all setups, extreme $\alpha$ values (near 0 or 1) underperform, and intermediate values (0.4 to 0.8) work best ; choosing an intermediate value for $\alpha$ helps to make a compromise between a clean graph topology and retrieving relevant graph entities. This supports the intuition of balancing the focus between subquestions and the main question during retrieval. In the rest of the paper and all following results, we use $\alpha=0.7$.